OpenStudy (anonymous):

square root of 40 plus square root of 250

OpenStudy (anonymous):

\[\sqrt{40}+\sqrt{250}=\sqrt{4*10}+\sqrt{25*10}=2\sqrt{10}+5\sqrt{10}=7\]

OpenStudy (anonymous):

sorry.\[7\sqrt{10}\]

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

yeah hang on she's correct.